OpenStudy (anonymous):

Your school drama club is putting on a play and hopes to raise $600 from ticket sales. They sell adult tickets for $8 each and student tickets for $5 each. Let x represent the number of adult tickets and y represent the number of student tickets. The equation to represent the number of tickets sold to reach their goal, 8x+ 5y = 600, is graphed below. If 30 student tickets are sold, how many adult tickets must be sold to reach their goal?

OpenStudy (anonymous):

A. 56.25 tickets B. 70 tickets C. 75 tickets D. 120 tickets

OpenStudy (anonymous):

Since y=student tickets, y=30. Plug this into the equation: 8x + 5(30) = 600 8x = 600 - 150 x = 450/8 x = 56.25

OpenStudy (anonymous):

thx